# AI Safety in the Eyes of the Downstream Developer: A First Look at Concerns, Practices, and Challenges

### v1

@@ 7. Conclusions
In this paper, we investigated the concerns, practices, and challenges downstream developers face in AI safety in AI-based software development. We conducted a mixed-method study by first interviewing 18 participants to derive qualitative insights for each RQ and triangulated the results by conducting a survey with 86 participants for quantitative perspectives. Additionally, we analysed incidents from an AI incident database to understand real-life AI safety failures to help assess developers’ awareness and approaches. Our findings reveal that despite high awareness of AI safety, developers take limited action during the initial preparation and PTM selection stages. Additionally, knowledge gaps, missing policies, and a lack of tool support contribute to widely varying approaches to AI safety. The frequent occurrence of AI incidents, particularly related to model reliability, harmful and biased content, and misuse and exploitation, raises concerns about the effectiveness of these approaches. We provide concrete implications for different stakeholders, emphasising the need to bridge the knowledge transfer gap between research/policy and downstream developers.
@@ 8. Data Availability
Data will be released upon acceptance.

### v3

@@ 7. Conclusions
In this paper, we investigated the concerns, practices, and challenges downstream developers face in AI safety in the development of AI-based software. We conducted a mixed-method study by first interviewing 16 participants to derive qualitative insights for each RQ and triangulated the results by conducting a survey with 86 participants for quantitative perspectives. Additionally, we analysed incidents from an AI incident database to understand real-life AI safety failures to help assess developers’ awareness and approaches.
The AI safety concerns reported by the downstream developers focus on data, model behaviour, and misuse/exploitation concerns. To address these concerns during development, downstream developers adopt several practices: 1) deeper understanding and assessment for safety concerns, 2) regular monitoring and assessment for safety concerns, 3) implementing technical safety guards, 4) maintaining documentation and expert consultation, and 5) relying on external parties and resources. The challenges they encounter include 1) infrastructure gaps, 2) technical and model-related difficulties, 3) lack of mature process and methodology, 4) resource constraints, 5) barriers to technical understanding and information access.
Our findings reveal that despite high awareness of AI safety, developers take limited action during the initial preparation and PTM selection stages. Additionally, knowledge gaps, missing policies, and a lack of tool support contribute to widely varying approaches to AI safety. The frequent occurrence of AI incidents, particularly related to model reliability, harmful and biased content, and misuse and exploitation, raises concerns about the effectiveness of these approaches. We provide concrete implications for different stakeholders, emphasising the need to bridge the knowledge transfer gap between research/policy and downstream developers.
@@ 8. Data Availability
Due to data privacy constraints, we cannot share raw interview transcripts. However, the interview protocol, coding process, survey screenshot, survey responses, and AI Incident Database analysis are available at https://github.com/Haoyu-Gao/AI-Safety-Mix-Study.